OpenStudy (anonymous):

If g(x) = x2 − 5, find g(7).

OpenStudy (anonymous):

plug 7 in for x

OpenStudy (bahrom7893):

yup so that u get: 7^2 - 5 = 49-5 = 44

OpenStudy (bahrom7893):

sorry lagrange but i think he meant square it..

OpenStudy (anonymous):

yeah yeah okay

OpenStudy (saifoo.khan):

Insert.